OpenStudy (anonymous):

Find the LCD for the fractions. -11/18 , 13/12

zepdrix (zepdrix):

Hey Savannah :) 18 = 3*6 12 = 2*6 So how many 6's do we need need to get a common denominator? 2*3 of them

zepdrix (zepdrix):

2*3*6 make sense? c:

OpenStudy (anonymous):

Yes, thank you! (: